# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### v2

@@ Abstract
We introduce SymbolicAI, a versatile and modular framework employing a logic-based approach to concept learning and flow management in generative processes. SymbolicAI enables the seamless integration of generative models with a diverse range of solvers by treating large language models (LLMs) as semantic parsers that execute tasks based on both natural and formal language instructions, thus bridging the gap between symbolic reasoning and generative AI. We leverage probabilistic programming principles to tackle complex tasks, and utilize differentiable and classical programming paradigms with their respective strengths. The framework introduces a set of polymorphic, compositional, and self-referential operations for data stream manipulation, aligning LLM outputs with user objectives. As a result, we can transition between the capabilities of various foundation models endowed with zero- and few-shot learning capabilities and specialized, fine-tuned models or solvers proficient in addressing specific problems. In turn, the framework facilitates the creation and evaluation of explainable computational graphs. We conclude by introducing a quality measure and its empirical score for evaluating these computational graphs, and propose a benchmark that compares various state-of-the-art LLMs across a set of complex workflows. We refer to the empirical score as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short. The framework codebase111 SymbolicAI framework released on January 20th, 2023, on GitHub: https://github.com/ExtensityAI/symbolicai. and benchmark222 Evaluation benchmark released on February 1st, 2024, on GitHub: https://github.com/ExtensityAI/benchmark. are linked below.
Figure 1: Our neuro-symbolic framework enables a seamless transition between classical and differentiable programming, each with distinct dynamics and strengths. Differentiable programming provides access to foundational and specialist models. Classical programming, on the other hand, shifts between abstraction and implementation, focusing on high-level concepts before delving into the details of implementation.
@@ 1 Introduction
The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022). These models have enhanced the functionality of tools for search-based interactions (YouWrite, 2022; Writesonic, 2022; Microsoft, 2023), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more. Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few. Therefore, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023). Yet, these applications merely scratch the surface of the transformation that language-based interactions are expected to bring to human-computer interactions in both the near and distant future.
In part, instruction-based fine-tuning of LLMs through reinforcement learning from human feedback (Ouyang et al., 2022; Li et al., 2023) or direct preference optimization (Rafailov et al., 2023) has shown promising results dealing with value misalignment issues (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Goyal et al., 2022), unlocking new possibilities for chain of thoughts (Wei et al., 2022b), tree of thoughts (Yao et al., 2023a), and graph of thoughts interactions (Besta et al., 2023). However, recent research also highlights the limitations of LLMs in functional linguistic competence despite their proficiency in formal linguistic competence (Mahowald et al., 2023). Whereas formal linguistic competence encompasses the ability to understand and generate language, functional linguistic competence pertains to the application of language in real-world contexts, such as conveying sensory input or recalling information from memory. Examples of functional linguistic competence include implicatures (Ruis et al., 2022) and contextual language comprehension beyond the statistical manifestation of data distributions (Bransford & Johnson, 1972). Consequently, operating LLMs through a purely inference-based approach confines their capabilities within their provided context window, severely limiting their horizon. This results in deficiencies for situational modeling, non-adaptability through contextual changes, and short-term problem-solving, amongst other capabilities. However, simply increasing the context length may not yield greater capabilities, as demonstrated by the observed U-shaped performance curve (Liu et al., 2023) where LLMs excel when using information at the beginning or end of the input context, but struggle with information located in the middle, especially as context increases. These challenges are actively being researched, with novel approaches such as Hyena (Poli et al., 2023), RWKV (Bo, 2021), GateLoop (Katsch, 2023), and Mamba (Gu & Dao, 2023) surfacing. Meanwhile, the re-emergence of interest in retrieval-augmented generative approaches (Li et al., 2022a) offers an alternative by circumventing the autoregressive nature of the widely-used Transformer architecture (Vaswani et al., 2017), enabling context enrichment with lateral information. In parallel, efforts have focused on developing tool-based approaches (Schick et al., 2023) or template frameworks (Chase, 2023) to extend large LLMs’ capabilities and enable a broader spectrum of applications. However, these efforts only partially capture the vast potential inherent in leveraging LLMs as semantic parsers.
@@
In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023). SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a). These operations guide the generative process and facilitate a modular design with a wide range of existing solvers, including formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval. It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers. In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016). We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
Lastly, alongside the framework, we introduce a benchmark and derive a quality measure and its empirical score to address the evaluation of multi-step NeSy generative processes. See Section 6 for more details.
In summary, we highlight the following key contributions of this work:
@@
  • We combine the broad applicability of LLMs as semantic parsers with symbolic expressions by leveraging a modular probabilistic programming paradigm, facilitating the creation of complex computational graphs.
  • We introduce a quality measure and its empirical score alongside a benchmark for comparing state-of-the-art LLMs across a wide range of tasks.
@@ 2 Related Work
The field of symbolic AI has its foundations in the works of the Logic Theorist (LT) (Newell & Simon, 1956) and the General Problem Solver (GPS) (Newell et al., 1957). These programs represented the first steps towards automated reasoning and problem-solving using symbolic representations. Despite their advancements, both faced challenges in dealing with the complexity of real-world problems, particularly due to the combinatorial nature of the solution space. To address these limitations, the Soar (Laird et al., 1987) cognitive architecture was developed, advancing the notion that intelligent behavior results from goal-oriented search through a problem space (Newell & Simon, 1972; McCarthy et al., 2006), with each step consisting of selecting and applying operators. Soar introduced components like reinforcement learning, impasses, substates, and chunking to enhance its problem-solving capabilities. It also demonstrated the importance of learning from experiences to adapt and improve performance over time. However, Santoro et al. (2022) emphasizes the subjectivity of symbols and suggests that human-like symbolic fluency could develop in machines through learning algorithms immersed in socio-cultural contexts. This perspective, anchored in the notion that symbols are triadic and their meaning emerges from consensus, seeks to move away from traditional symbolic AI methodologies towards AI that adaptively learns meaning and behaviors from human-like experiences. The goal is to cultivate machines that demonstrate symbolic behaviors across a spectrum of competencies, potentially mirroring the evolutionary and social learning processes observed in humans. Lastly, symbolic AI struggles with real-world data’s unpredictability and variability. These challenges have led to the employment of statistical learning methodologies, like deep learning (Alom et al., 2018), which are more adept at managing noise and uncertain information through vector-valued representations.
@@
The sub-symbolic framework, rooted in neural network paradigms, began with pioneering works such as the perceptron (McCulloch & Pitts, 1943), with the first hardware implementation quickly following (Rosenblatt, 1958). The foundational notion of distributed processing (Rumelhart et al., 1986) was later bolstered and further expanded by demonstrating that multilayer feedforward networks with a single hidden layer can serve as universal approximators for any Borel measurable function, given sufficient hidden units (Hornik et al., 1989). Fast-forward, contemporary frameworks achieve a significant leap with the introduction of the Transformer architecture (Vaswani et al., 2017), which underpins most of today’s LLMs. These LLMs demonstrate exceptional capabilities in in-context learning, a method popularized by the likes of GPT-3 (Brown et al., 2020), where models improve task performance through natural language instruction and examples provided directly in the input prompt. While in-context learning bypasses the need for explicit retraining, it demands meticulous prompt design to steer models towards desired behaviors. Despite their versatility, current LLMs face challenges such as fallacious reasoning and the generation of erroneous content, commonly referred to as hallucinations (Jones & Steinhardt, 2022). These limitations highlight the importance of integrating complementary symbolic methods to validate and guide the generative processes of LLMs, ensuring more accurate and reliable outputs.
@@
To overcome the limitations of each individual method, NeSy approaches meld the statistical inference strengths of deep neural architectures with the generalization and explainability of symbolic systems (Besold et al., 2017; Yu et al., 2023; Hamilton et al., 2022; Garcez et al., 2015; d’Avila Garcez et al., 2019; d’Avila Garcez & Lamb, 2020; Lamb et al., 2020). Some approaches focus on different strategies for integrating learning and reasoning processes (Yu et al., 2023; Fang et al., 2024). Firstly, learning for reasoning methods treat the learning aspect as an accelerator for reasoning, in which deep neural networks are employed to reduce the search space for symbolic systems (Qu & Tang, 2019; Silver et al., 2016, 2017b, 2017a; Schrittwieser et al., 2020). Secondly, reasoning for learning views reasoning as a way to regularize learning, in which symbolic knowledge acts as a guiding constraint that oversees machine learning tasks (Hu et al., 2016; Xu et al., 2018). Thirdly, the learning-reasoning category enables a symbiotic relationship between learning and reasoning. Here, both elements interact and share information to boost problem-solving capabilities (Donadello et al., 2017; Manhaeve et al., 2018; Mao et al., 2019; Ellis, 2023). This synergy further extends when considering graph-based methods, which closely align with the objectives of our proposed framework. Research in this area, such as CycleGT (Guo et al., 2020) and Paper2vec (Ganguly & Pudi, 2017), explored unsupervised techniques for bridging graph and text representations. Subsequently, graph embeddings, when utilized within symbolic frameworks, can enhance knowledge graph reasoning tasks (Zhang et al., 2021), or more generally, provide the bedrock for learning domain-invariant representations (Park et al., 2023). Lastly, building upon the insights from Sun et al. (2022), the integration of NeSy techniques in scientific workflows promises significant acceleration in scientific discovery. While previous work has effectively identified opportunities and challenges, we have taken a more ambitious approach by developing a comprehensive framework from the ground up to facilitate a wide range of NeSy integrations.

### v3

@@ Abstract
We introduce SymbolicAI, a versatile and modular framework employing a logic-based approach to concept learning and flow management in generative processes. SymbolicAI enables the seamless integration of generative models with a diverse range of solvers by treating large language models (LLMs) as semantic parsers that execute tasks based on both natural and formal language instructions, thus bridging the gap between symbolic reasoning and generative AI. We leverage probabilistic programming principles to tackle complex tasks, and utilize differentiable and classical programming paradigms with their respective strengths. The framework introduces a set of polymorphic, compositional, and self-referential operations for multi-modal data that connects multi-step generative processes and aligns their outputs with user objectives in complex workflows. As a result, we can transition between the capabilities of various foundation models with in-context learning capabilities and specialized, fine-tuned models or solvers proficient in addressing specific problems. Through these operations based on in-context learning our framework enables the creation and evaluation of explainable computational graphs. Finally, we introduce a quality measure and its empirical score for evaluating these computational graphs, and propose a benchmark that compares various state-of-the-art LLMs across a set of complex workflows. We refer to the empirical score as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short. The framework codebase111 SymbolicAI framework: https://github.com/ExtensityAI/symbolicai and benchmark222 Evaluation benchmark: https://github.com/ExtensityAI/benchmark
     * Correspondence to: dinu@ml.jku.at, {marius, leo}@extensity.ai
Work done outside of Amazon.
are linked below.
Figure 1: Our neuro-symbolic framework enables a seamless transition between symbolic and differentiable programming, each with distinct dynamics and strengths. Differentiable programming provides access to foundational and specialist models. Classical programming, on the other hand, shifts between abstraction and implementation, focusing on high-level concepts before delving into the details of implementation.
@@ 1 Introduction
The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022). These models have enhanced the functionality of tools for search-based interactions (YouWrite, 2022; Writesonic, 2022; Microsoft, 2023), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more. Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few. Consequently, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023).
Despite their versatility, current LLMs face challenges such as fallacious reasoning and the generation of erroneous content, commonly referred to as hallucinations (Jones & Steinhardt, 2022). These limitations highlight the importance of integrating complementary symbolic methods to validate and guide the generative processes of LLMs, ensuring more accurate and reliable outputs. In parallel, efforts have focused on developing tool-based approaches (Schick et al., 2023) or template frameworks (Chase, 2023) to extend LLMs’ capabilities and enable a broader spectrum of applications. However, these efforts only partially capture the potential inherent in leveraging LLMs as semantic parsers. In contrast to parsers for structured languages a semantic parser is able to break down unstructured human language into semantically meaningful components and transform those into a structured form. While traditionally semantic parsing has been a role filled by specialized algorithms and models, we posit that LLMs, through their training on diverse linguistic data, have developed the ability to perform semantic parsing as part of their broader natural language processing capabilities. In turn, we identify LLMs as a central component in creating sophisticated neuro-symbolic (NeSy) AI systems. These systems integrate symbolic and sub-symbolic concepts and utilize the capabilities of semantic parsing to develop symbolic expressions that enable new probabilistic programming paradigms.
@@
We introduce SymbolicAI, a compositional NeSy framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023). SymbolicAI augments the generative process of LLMs with in-context learning operations, realized through functional primitives, and enables the creation of versatile applications through in-context learning (Wei et al., 2022a). These operations enable logic-based components that guide the generative process and enable a modular NeSy system, including a wide range of existing solvers, formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval. SymbolicAI exposes these solvers as building blocks for constructing compositional functions as computational graphs, making it possible to bridge classical and differentiable programming paradigms with the aim to create domain-invariant problem solvers. In designing the architecture of SymbolicAI, we drew inspiration from a body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016). We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022; 2023). We hypothesize that such a central processing module based in language is a core component of broad AI systems (see Appendix Section A) and enables the development of fully autonomous AI systems for decision-making.
A significant challenge encountered in the development of our framework pertained to the evaluation of LLMs when used as semantic parsers in a NeSy workflow. Current evaluation of generated content relies on metrics for single-step generative processes, such as the BLEU score (Papineni et al., 2002). These metrics are not suitable for evaluating multi-step generative processes. BLEU has limitations, as it measures $n$-gram-based overlap of generated output with a reference that does not consider the semantic meaning. As a result, BLEU fails to capture semantic equivalence, especially in more complex tasks. More recent metrics such as CIDEr (Vedantam et al., 2014) or SPICE (Anderson et al., 2016) are also not suitable for our problem, either because they are built upon BLEU (in case of CIDEr) or designed with inductive biases specifically for image captioning.
Therefore, alongside our framework we introduce a quality measure (and its empirical score) for multi-step generative processes based on semantic meaning. We refer to our score as ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short. Our VERTEX score uses embeddings to compare node distributions within a computational graph. It measures the semantic meaning across the distributional path by computing at each node the cross-similarity between the generated embeddings and embeddings sampled from a reference distribution. Furthermore, the VERTEX score is designed such that it can be used as a reward signal in a reinforcement learning setting (Sutton, 1984). Finally, we propose a benchmark for evaluating complex workflows. We define a set of basic evaluations, particularly associative predictions based on in-context learning, multi-modal bindings for tool utilization, and program synthesis for subroutine execution. Furthermore, we introduce complex evaluations for logic-based components and hierarchical computational graphs.
In summary, the key contributions presented in this work are as follows:
@@
  • We leverage LLMs as semantic parsers to enable the creation of complex computational graphs by combining symbolic expressions with probabilistic programming paradigms.
  • We introduce a quality measure and its empirical score alongside a benchmark designed for multi-step generative processes for comparing LLMs across a wide range of complex tasks.
@@ 2 Related Work
The field of symbolic AI has its foundations in the works of the Logic Theorist (LT) (Newell & Simon, 1956) and the General Problem Solver (GPS) (Newell et al., 1957). These programs represented the first steps towards automated reasoning and problem-solving utilizing symbolic representations. Despite their advancements, both faced challenges in dealing with the complexity of real-world problems, particularly due to the combinatorial nature of the solution space. To address these limitations, the Soar (Laird et al., 1987) cognitive architecture was developed, advancing the notion that intelligent behavior results from goal-oriented search through a problem space (Newell & Simon, 1972; McCarthy et al., 2006), with each step consisting of selecting and applying operators. Soar introduced components like reinforcement learning, impasses, sub-states, and chunking to enhance its problem-solving capabilities. It also demonstrated the importance of learning from experiences to adapt and improve performance over time. However, Santoro et al. (2022) emphasizes the subjectivity of symbols and suggests that human-like symbolic fluency could develop in machines through learning algorithms immersed in socio-cultural contexts. This perspective, anchored in the notion that symbols are triadic and their meaning emerges from consensus, seeks to move away from traditional symbolic AI methodologies towards AI that adaptively learns meaning and behaviors from human-like experiences. The goal is to cultivate machines that demonstrate symbolic behaviors across a spectrum of competencies, potentially mirroring the evolutionary and social learning processes observed in humans. Lastly, symbolic AI struggles with real-world data’s unpredictability and variability. These challenges have led to the employment of statistical learning methodologies, like deep learning (Alom et al., 2018), which are more adept at managing noise and uncertain information through vector-valued representations.
@@
The sub-symbolic framework, rooted in neural network paradigms, began with pioneering works such as the perceptron (McCulloch & Pitts, 1943), with the first hardware implementation quickly following (Rosenblatt, 1958). The foundational notion of distributed processing (Rumelhart et al., 1986) was later bolstered and further expanded by demonstrating that multilayer feedforward networks with a single hidden layer can serve as universal approximators for any Borel measurable function, given sufficient hidden units (Hornik et al., 1989). Fast-forward, contemporary frameworks achieve a significant leap with the introduction of the Transformer architecture (Vaswani et al., 2017), which underpins most of today’s LLMs. These LLMs demonstrate exceptional capabilities in in-context learning, a method popularized by the likes of GPT-3 (Brown et al., 2020), where models improve task performance through natural language instruction and examples provided directly in the input prompt. While in-context learning bypasses the need for explicit retraining, it demands meticulous prompt design to steer models towards desired behaviors.
@@
To overcome the limitations of each individual method, NeSy approaches meld the statistical inference strengths of deep neural architectures with the generalization and explainability of symbolic systems (Garcez et al., 2015; Besold et al., 2017; d’Avila Garcez et al., 2019; d’Avila Garcez & Lamb, 2020; Lamb et al., 2020; Hamilton et al., 2022; Yu et al., 2023). Some approaches focus on different strategies for integrating learning and reasoning processes (Yu et al., 2023; Fang et al., 2024). Firstly, learning for reasoning methods treat the learning aspect as an accelerator for reasoning, in which deep neural networks are employed to reduce the search space for symbolic systems (Silver et al., 2016; 2017b; 2017a; Qu & Tang, 2019; Schrittwieser et al., 2020). Secondly, reasoning for learning views reasoning as a way to regularize learning, in which symbolic knowledge acts as a guiding constraint that oversees machine learning tasks (Hu et al., 2016; Xu et al., 2018). Thirdly, the learning-reasoning category enables a symbiotic relationship between learning and reasoning. Here, both elements interact and share information to boost problem-solving capabilities (Donadello et al., 2017; Manhaeve et al., 2018; Mao et al., 2019; Ellis, 2023). This synergy further extends when considering graph-based methods, which closely align with the objectives of our proposed framework. Research in this area, such as CycleGT (Guo et al., 2020) and Paper2vec (Ganguly & Pudi, 2017) explores unsupervised techniques for bridging graph and text representations, GPTSwarm (Zhuge et al., 2024) explores graph optimizers to refine node-level prompts and edge optimization. Subsequently, graph embeddings, when utilized within symbolic frameworks, can enhance knowledge graph reasoning tasks (Zhang et al., 2021), or more generally, provide the bedrock for learning domain-invariant representations (Park et al., 2023).
Lastly, building upon the insights from Sun et al. (2022), the integration of NeSy techniques in scientific workflows promises significant acceleration in scientific discovery. While previous work has effectively identified opportunities and challenges, we have taken a more ambitious approach by developing a comprehensive framework from the ground up to facilitate a wide range of NeSy integrations.
Large Language Models
In part, instruction-based fine-tuning of LLMs through reinforcement learning from human feedback (Ouyang et al., 2022; Li et al., 2023) or direct preference optimization (Rafailov et al., 2023) has shown promising results dealing with value misalignment issues (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Goyal et al., 2022), unlocking new possibilities for chain of thoughts (Wei et al., 2022b), tree of thoughts (Yao et al., 2023a), and graph of thoughts interactions (Besta et al., 2023). However, recent research also highlights the limitations of LLMs in functional linguistic competence despite their proficiency in formal linguistic competence (Mahowald et al., 2023). Whereas formal linguistic competence encompasses the ability to understand and generate language, functional linguistic competence pertains to the application of language in real-world contexts, such as conveying sensory input or recalling information from memory. Examples of functional linguistic competence include implicatures (Ruis et al., 2022) and contextual language comprehension beyond the statistical manifestation of data distributions (Bransford & Johnson, 1972; Mikolov et al., 2013b). Consequently, operating LLMs through a purely inference-based approach confines their capabilities within their provided context window, severely limiting their horizon. This results in deficiencies for situational modeling, non-adaptability through contextual changes, and short-term problem-solving, amongst other capabilities. However, simply increasing the context length may not yield greater capabilities, as demonstrated by the observed U-shaped performance curve (Liu et al., 2023) where LLMs excel when utilizing information at the beginning or end of the input context, but struggle with information located in the middle, especially as context increases. These challenges are actively being researched, with novel approaches such as Hyena (Poli et al., 2023), RWKV (Bo, 2021), GateLoop (Katsch, 2023), Mamba (Gu & Dao, 2023) and xLSTM (Beck et al., 2024) surfacing. Meanwhile, the re-emergence of interest in retrieval-augmented generative approaches (Li et al., 2022a) offers an alternative by circumventing the autoregressive nature of the widely-utilized Transformer architecture (Vaswani et al., 2017), enabling context enrichment with lateral information.
In-Context Learning
Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023). This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Wei et al., 2022b; Singhal et al., 2023). CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output. Self-generated CoT, where models are encouraged to generate their own reasoning chains based on training examples, surpasses even expertly crafted CoT (Fernando et al., 2023). This observation echoes other reports that GPT-4 has an emergent self-improving capability through introspection, such as self-verification (Weng et al., 2023) or self-consistency (Wang et al., 2023b). Tree of Thoughts (ToT) enables LLMs to solve complex problems by exploring multiple reasoning paths through a search tree of coherent text units, demonstrating significant problem-solving enhancements in tasks requiring strategic planning and search (Yao et al., 2023a). Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023).